A car gets 18 MPG in city driving and 24 MPG in highway driving. The car is driven 465 miles on 23 gallons of gasoline. How may miles were driven in the city and how many miles were driven on the highway?
Answer by stanbon(75887)

A car gets 18 MPG in city driving and 24 MPG in highway driving. The car is driven 465 miles on 23 gallons of gasoline. How may miles were driven in the city and how many miles were driven on the highway?
--------------------
Equation:
miles + miles = miles
18x + 24(23-x) = 465 miles
18x + 24*23 - 24x = 465
-6x = 465-24*23
-6x = -87
x = 14.5 gallons
---------------------------
highway distance = 24(23-14.5) = 204 miles
===============================================
Cheers,
Stan H.